OpenStudy (jaweria):

Help me Please!! y=cos^-1(sin x+5).

OpenStudy (jaweria):

\[Y=\cos^{-1} (\sin x+5).\]

OpenStudy (saifoo.khan):

And what do we have to do with this equation?

OpenStudy (jaweria):

We have to take derivative

OpenStudy (saifoo.khan):

@satellite73

OpenStudy (anonymous):

there is no such thing!

OpenStudy (anonymous):

\[4\leq \sin(x)+5\leq 6\] and that is not in the domain of arccosine

OpenStudy (jaweria):

what is that?

OpenStudy (anonymous):

the derivative of vos^-1(x) \[\Large \frac{d}{dx}(\cos^{-1}(x))=\frac{-1}{\sqrt{1-x^2}}\] replace x with sinx+5 use chanin rule \[\Large \frac{d}{dx}(\cos^-1{(\sin x+5)}=\frac{-1}{\sqrt{1-(\sin(x)+5)^2}}*\frac{d}{dx}(\sin(x+5))\]

OpenStudy (jaweria):

but what will be my answer for this equation? can you please help all the way till the answer?

OpenStudy (anonymous):

last time i provided answer was reported . try it i will guide you .

OpenStudy (anonymous):

maybe it is \[\cos^{-1}(\sin(x+5))\]

OpenStudy (anonymous):

if so, then start with \[\cos^{-1}(\sin(x+5))=\sqrt{1-(x+5)^2}\] and differentiate that